Find all combinations of truth values for p, q and r for which the statement ¬p ↔ (q ∧ ¬(p → r)) is true


1
Expert's answer
2021-04-19T01:40:08-0400

Let's compose a truth table:



Then  the statement ¬p ↔ (q ∧ ¬(p → r)) is true if"\left( {p = True,\,\,q = r = False} \right);\,\,\,\left( {p = r = True,q = False} \right);\,\,\left( {p = q = r = True} \right)"

Answer: "\left( {p = True,\,\,q = r = False} \right);\,\,\,\left( {p = r = True,q = False} \right);\,\,\left( {p = q = r = True} \right)"